Find both the x-intercept and the y-intercept of the line given equation:
5.6x-3.8y+9.6=0

Answer by Alan3354(69443)

Find both the x-intercept and the y-intercept of the line given equation:
5.6x-3.8y+9.6=0
-------
5.6x-3.8y = -9.6
At the x-int, y=0
5.6x = -9.6
x = -9.6/5.6
x = - 12/7
==============
At the y-int, x=0
etc